OpenStudy (anonymous):

can someone help me, washers/ disk method of y^3 = x, y=0, x=1 about the y-axis?

OpenStudy (anonymous):

first write the equation in terms of y= ....

OpenStudy (anonymous):

y=x^3, I keep getting pi over seven and its 6pi over 7

OpenStudy (anonymous):

did you change your equation in the question?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

i thought the equation originally said y^(1/3)=x but know it says y^3=x

OpenStudy (anonymous):

i could have just saw in wrong sorry

OpenStudy (anonymous):

i worked it out too and got pi/6

OpenStudy (anonymous):

can you show me?

OpenStudy (anonymous):

it's like pi(y^3)^2 and then i integrate and its 1/7 y^7 and i plug in between 1 and 0 and get pi over seven ?

OpenStudy (anonymous):

yep thats what i did

OpenStudy (anonymous):

where does the six come from?

OpenStudy (anonymous):

oh I'm sorry, thats supposed to be a 7

OpenStudy (anonymous):

the answer according to the shell method is 6pi over seven.

OpenStudy (anonymous):

so for the shell method integral of (2*pi*x*x^1/3)dx = 2*pi*x^(7/3)*(3/7)=6pi/7

OpenStudy (anonymous):

if you use different methods are you supposed to get different answers?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

could you show me the shell way of solving it, to get 6 pi over 7?

OpenStudy (anonymous):

integral of (2*pi*x*x^1/3)dx = 2*pi*x^(7/3)*(3/7)=6pi/7

OpenStudy (anonymous):

do you need it broken down more

OpenStudy (anonymous):

sorry i meant the washer way.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

just by the washer method because when oh okay

OpenStudy (anonymous):

well when your adding up washers the volume is pi*r^2*h for the height its infinityls thin so h = dy. the radius is your distance from the x axis so y^3. your going from y=0 to y=1 so its integral of (pi*(y^3)^2*dy)= pi*y^7/7=pi/7

OpenStudy (anonymous):

thats what i get at least

OpenStudy (anonymous):

but the answer is 6 pi over 7, you don't know why there is a different answer?

OpenStudy (anonymous):

i really don't. im not sure, but i believe im doing the method correct.

OpenStudy (anonymous):

me too, okay well thanks for giving it your best attempt. i'm stymied.

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

wait a second,

OpenStudy (anonymous):

if you look at the graph of the function you'll find when you are doing the washer methed you get the volume whats above the curve, but the shell method will give you the volume under the curve

OpenStudy (anonymous):

so they will be different answers

OpenStudy (anonymous):

at least the way we set them up.